Tyler has been saving his winning lottery tickets. He has 23 tickets that are worth a total of $175. If each ticket is worth either $5 or $10, how many of each does he have?
Answer by fractalier(6550)

Call the numbers of $5 and $10 tickets, x and y. Then we write
x + y = 23 and the value equation is
5x + 10y = 175
Multiply the top one by 5 and subtract...we get
5x + 10y = 175
-(5x + 5y = 115)
-------------------
5y = 60
y = 12 $10 tickets
which means there are
x = 11 $5 tickets